OpenStudy (anonymous):

How does sqrt(3k^3) simplify? does it become (k)sqrt(3)

OpenStudy (dumbcow):

if it is a cubed root, then yes. If its a square root, then no

OpenStudy (anonymous):

How does it behave if it is a square root?

OpenStudy (dumbcow):

sqrt(k^3) or k^(3/2)

OpenStudy (anonymous):

sqrt(k^3)

OpenStudy (anonymous):

\[\sqrt{3k ^{2}}=\sqrt{3k.k ^{2}}=k \sqrt{3k}\]